# MESA-QUEST: Tracing the formation of direct collapse black hole seeds via quasi-stars

### v2

@@ 4 Previous Work on quasi-stars
Quasi-stars were first proposed by M. C. Begelman et al. (2008) where they modeled these objects analytically using loaded polytropes. It was suggested that quasi-stars could exist long enough for the central black hole to accrete enough material to form a heavy seed with masses $\sim 10^{4-5}$ M by $z\gtrsim 10$. The unique characteristic of these objects, and what makes them appealing as heavy-seed progenitors, is that the black hole is able to accrete at the Eddington limit of the entire object, not just the black hole itself. This extreme accretion rate is due to the formation of an accretion disk very near the black hole, as well as the convection that occurs just outside it. This then allows for much more efficient release of energy and an increase in the accretion rate of the black hole. M. C. Begelman et al. (2006) suggested that after reaching a certain photospheric temperature, the quasi-star may eventually eject its envelope, imposing a limit on the attainable black hole mass based on the mass of the entire object.
@@
W. H. Ball (2012) numerically simulated quasi-stars using the Cambridge STARS code with a spherically symmetric Bondi-Hoyle accretion scheme and inner boundary condition. They found that there was a stronger limit on black hole growth dependent on the black hole to total mass ratio than suggested by M. C. Begelman et al. (2006), showing that this ratio could not exceed $\sim 0.167$. They argue that this limit is similar to that of the Schönberg-Chandrasekhar limit, though the exact physical explanation remains to be determined.
@@
E. R. Coughlin & M. C. Begelman (2024) explored this limit further, suggesting that it is due to a non-physical formulation of the inner boundary condition. They modeled a deeper, maximally convective region well within the previous boundary, where energy is transported outward by strong convection. Their “saturated-convection” boundary condition leads to a black hole to total mass ratio of 0.6~0.6.
@@
Quasi-stars must remain stable long enough for the central black hole to accrete a sufficient amount of material to create a heavy seed. There are then a number of stability concerns when it comes to this sustainability, including the effects of surface winds. C. Dotan et al. (2011) argued that radiation-driven winds are prominent enough to unravel the envelope on a shorter timescale than that of black hole accretion, stifling the black hole growth for all quasi-star masses below a few $\times~10^{5}$ M. W. H. Ball (2012) employed the Reimers wind scheme, a simple analytic wind based on the mass, radius, and luminosity of the star (D. Reimers, 1975) and show that the winds have very little effect on the lifetime of the star and thus the final black hole mass. However, this scheme is likely not appropriate for the extreme conditions in supermassive or quasi-stars and was implemented as a first approximation. D. Fiacconi & E. M. Rossi (2015) suggested that taking into account a previously neglected advection term in the radiation-driven wind calculation used by C. Dotan et al. (2011) significantly decreases the impact of the wind on the evaporation timescale of the quasi-star.
5 MESA Implementation of Quasi-Star evolution
@@
Modules for Experiments in Stellar Astrophysics (MESA) (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) is a collection of open-source suites designed to tackle a wide range of problems in computational stellar astrophysics. In this work, we are particularly interested in the 1-D stellar evolution module, MESA star. This module utilizes adaptive mesh refinement and advanced timestep controls to solve fully coupled structure and composition equations. MESA also uses a series of independently constructed modules to provide equation of state, opacity, nuclear reaction rates, element diffusion date, and atmosphere boundary conditions.
5.1 Quasi-stars in MESA
@@
In previous works, we developed modifications to MESA in order to model the quasi-star scenario (C. B. Campbell et al., 2025). The black hole was implemented through adjustment of the boundary conditions and by including a sub-grid model for the black hole accretion and luminosity. Due to the massive black hole at the center of a quasi-star, general relativistic effects become important. We account for this using the Tolman-Oppenheimer-Volkov correction – a modification to the equation for hydrostatic equilibrium via a change to the gravitational constant $G$, which is recalculated within MESA at every timestep. Additionally, we use the Ledoux criterion for convection along with a mixing length of $\alpha_{mlt}=2$ as has also been used in previous models for supermassive stars (B. Paxton et al., 2011; N. P. Herrington et al., 2023).
@@
To create an initial model, we evolve a typical stellar model with an initial mass of 2020 M and an envelope accretion rate of $0.1~\textrm{M}_{\odot}\textrm{yr}^{-1}$ until it reaches $10,000$ M, which is the total object mass for our fiducial quasi-star models. This method is similar to those used in N. P. Herrington et al. (2023) when modeling supermassive stars. From this initial model, we then implement the aforementioned methods to account for the central black hole. This approach to initialization is not always necessary but does not significantly alter the quasi-star structure or lifetime in an impactful way and can thus be used regardless of necessity.
@@ 5.2 Boundary Conditions
where $\rho_{0}$ is the density of the innermost cell, and $n=1$ or $3$ according to the structure of the surrounding star and whether the angular momentum is transported inward or outward, respectively (E. Quataert & A. Gruzinov, 2000). We then solve the integral
@@
5.2.1 Bondi Radius
The first inner boundary condition we adopt is the same as originally used by W. H. Ball (2012) – the Bondi radius $R_{B}$. This is the radius at which the thermal energy of the particles equals that of the gravitational potential energy such that:
@@ 5.2.1 Bondi Radius
where $m$ is the particle mass, $c_{s}$ is the sound speed, $M_{\mathrm{BH}}$ is the black hole mass and $G$ is the gravitational constant. Anything within this radius will eventually fall into the black hole, although local fluid motions within may still move both toward and away from the black hole (W. H. Ball, 2012; W. H. Ball et al., 2012).
5.2.2 Saturated-Convection Radius
Recently E. R. Coughlin & M. C. Begelman (2024) proposed a new, tighter inner boundary condition for quasi-stars where convection transports energy outward at the local maximum efficiency. The result is a deeper hydrostatic region than obtained with the original Bondi conditions because the material drifts slowly inwards, compared to the free fall that the Bondi condition assumes.
@@ 5.2.2 Saturated-Convection Radius
These equations use dimensionless parameters $\xi=r/r_{\rm i}$ and $m_{\rm i}=m(r)/M_{\rm i}$, where $M_{\rm i}$ is the mass contained within $r_{\rm i}$. In this case, $K_{i}$ is the only free parameter to solve saturated convection solutions, and therefore can be used to determine the inner radius given the luminosity and mass of the object (E. R. Coughlin & M. C. Begelman, 2024). Additionally, our density profile must now be $\rho\propto r^{-1/2}$ if convection is to transport all of the energy around the black hole, similar to a convection dominated accretion flow (CDAF) solution E. Quataert & A. Gruzinov (2000). We then adjust $M_{\mathrm{cav}}$ accordingly.
5.3 Accretion & Luminosity
We now discuss the accretion schemes we have implemented within MESA for modeling quasi-stars. The defining characteristic of quasi-stars, and what makes them a viable heavy seed formation channel, is the fact that they are able to accrete at the Eddington rate of the entire object (M. C. Begelman et al., 2006). This allows the seed to grow within the appropriate time constraints as set by the observation of the earliest SMBHs (M. Castellano et al., 2022; R. P. Naidu et al., 2022; Y. Harikane et al., 2023; I. Labbé et al., 2023; P. Natarajan et al., 2024).
We first begin with the convection-limited Bondi accretion rate, discussed previously in W. H. Ball (2012). This is given by:
@@ 5.3 Accretion & Luminosity
Using this accretion rate and luminosity along with the Bondi radius, we obtain qualitatively similar results to those found in Part 3 of W. H. Ball (2012) (C. B. Campbell et al., 2025).
@@
Our fiducial models, shown in Fig. 1, begin with a central black hole mass of 10 M and a total mass of $10^{4}$ M. We used a primordial composition of 75% hydrogen and 25% helium and chose $\alpha$ values of 0.5, 0.8, 1.0, and 1.4 to explore both potential regimes where the accretion rate is suppressed (e.g., by angular momentum) and where it is enhanced (e.g., by magnetic fields).
5.4 Results
@@
We observe similar final black hole masses to those seen in W. H. Ball et al. (2012) and E. R. Coughlin & M. C. Begelman (2024) for the Bondi radius and saturated-convection radius respectively, with a factor of $\sim 5$ increase in the saturated-convection radius over the Bondi radius for $\alpha=1$ and $1.4$. Our models then also show that these mass limits are indeed not physical, but rather a limitation of the integration scheme – the entire calculation, and thus the black hole growth, halts when the inner boundary reaches the surface of the quasi-star.
@@
Of note is the step-like increase in the Bondi radius. This is likely due to the connection of convective regions that were previously separated by interstitial radiative regions (W. H. Ball, 2012). This merging would cause the star to become fully convective, sharply decreasing the sound speed, thus increasing the Bondi radius.
@@
6 Winds & Mass Loss
Up to this point, our models have only included mass loss due to the conversion of a portion of the accreted mass to photons. However, mass loss due to surface winds is far more prominent and may occur at a much higher rate (C. Dotan et al., 2011). Winds in “typical” stars are fairly well understood and have a wide variety of models for various star types (D. Reimers, 1975; C. de Jager et al., 1988; H. Nieuwenhuijzen & C. de Jager, 1990; T. Bloecker, 1995; T. Nugis & H. J. G. L. M. Lamers, 2000; J. S. Vink et al., 2001; J. T. van Loon et al., 2005; R. Björklund et al., 2021). In this section, we explore a few of these models that are built into MESA and that best satisfy the conditions within quasi-stars as first-principle approximations. Additionally, we implement a super-Eddington radiation-driven wind scheme from C. Dotan et al. (2011) that is specific to supermassive and quasi-stars.
6.1 Reimers Scheme
Our first wind scheme is that of D. Reimers (1975), which was also implemented in W. H. Ball (2012) and will be referred to hereafter as the Reimers scheme. This scheme is commonly used for red giants and includes a scaling factor $\eta$ to allow for different mass loss efficiencies. The mass loss can be analytically calculated and is given by:
@@ 6.1 Reimers Scheme
where $M$, $L$, and $R$ are the total mass, luminosity, and radius in solar units, respectively, and we use a typical value $\eta=0.5$ (D. Reimers, 1975; B. Paxton et al., 2011).
6.2 “Dutch” Scheme
Next we choose to explore a wind scheme used frequently for massive stars, defined as the “Dutch” scheme within MESA. This scheme is a combination of the wind models from C. de Jager et al. (1988), J. T. van Loon et al. (2005), and T. Nugis & H. J. G. L. M. Lamers (2000) where the transitions between them occur based on the surface temperature. Specifically, the combination used is based on E. Glebbeek et al. (2009) where they formulated the scheme to work for stars with masses $a~few\times 100$ M. For surface temperature $T_{\mathrm{eff}}>10^{4}$ K, the scheme of J. T. van Loon et al. (2005) is used while that of C. de Jager et al. (1988) is used for $T_{\mathrm{eff}}\leq 10^{4}$ K. It should be noted that the third model in the Dutch scheme of T. Nugis & H. J. G. L. M. Lamers (2000) is not used for our purposes, since it only occurs when the hydrogen mass fraction is <0.4<0.4. The temperatures and luminosities involved in the massive stars that the Dutch scheme was formulated for make this a better fit for quasi-stars, though it is likely still not appropriate. However, this scheme is built into MESA and serves as an improvement on the Reimers scheme.
@@
6.3 Super-Eddington Radiation-Driven Wind Schemes
Our final wind schemes come from C. Dotan et al. (2011) and D. Fiacconi & E. M. Rossi (2015) which were derived specifically for quasi-stars. These schemes account for the effects of the super-Eddington luminosities and the accompanying outflows. C. Dotan et al. (2011) calculates approximate hydrostatic solutions for the accretion powered envelope allowing for mass loss at the surface, resulting in a mass loss rate given by:
@@ 6.3 Super-Eddington Radiation-Driven Wind Schemes
where $R_{\star}$, $\rho_{\star}$, and $c_{\star}$ are the radius, density, and sound speed at the surface, respectively. The addition of the Mach number $\mu$ allows the case where the sonic point (i.e., where the gas velocity is $c_{s}$) is not the critical point where the net forces balance each other (i.e., $R_{\star}$ when in equilibrium) (C. Dotan et al., 2011; D. Fiacconi & E. M. Rossi, 2015).
This scheme, however, was not modeled solving the full equations of motion and neglected the term for advection energy. D. Fiacconi & E. M. Rossi (2015) continued to numerically solve these equations with the inclusion of both diffusion and advection of energy. They found that the mass loss rate due to these winds can be reasonably well fit with power laws in the $M_{\textrm{BH}}$-M plane using:
@@
We implement this as our radiation-driven wind scheme within MESA, with a few caveats. Firstly, this model does not take into account acceleration of the material when leaving the envelope. This acceleration occurs at a radius on the length-scale of the density scale height. However, on average this scale-height is of order $1\%~R_{\star}$, so can safely be ignored for this work. This scheme also does not account for the effects of rotation, which is outside the scope of this work (D. Fiacconi & E. M. Rossi, 2015).
6.4 Results

### v3

@@ IV Previous Work on quasi-stars
Quasi-stars were first proposed by Begelman_2008 where they modeled these objects analytically using loaded polytropes. It was suggested that quasi-stars could exist long enough for the central black hole to accrete enough material to form a heavy seed with masses $\sim 10^{4-5}$ M by $z\gtrsim 10$. The unique characteristic of these objects, and what makes them appealing as heavy-seed progenitors, is that the black hole is able to accrete at the Eddington limit of the entire object, not just the black hole itself. This extreme accretion rate is due to the formation of an accretion disk very near the black hole, as well as the convection that occurs just outside it. This then allows for much more efficient release of energy and an increase in the accretion rate of the black hole. begelman2006 suggested that after reaching a certain photospheric temperature, the quasi-star may eventually eject its envelope, imposing a limit on the attainable black hole mass based on the mass of the entire object.
@@
ball2012quasistars numerically simulated quasi-stars using the Cambridge STARS code with a spherically symmetric Bondi-Hoyle accretion scheme and inner boundary condition. They found that there was a stronger limit on black hole growth dependent on the black hole to total mass ratio than suggested by begelman2006, showing that this ratio could not exceed $\sim 0.167$. They argue that this limit is similar to that of the Schönberg-Chandrasekhar limit, though the exact physical explanation remains to be determined.
@@
coughlin2024 explored this limit further, suggesting that it is due to a non-physical formulation of the inner boundary condition. They modeled a deeper, maximally convective region well within the previous boundary, where energy is transported outward by strong convection. Their “saturated-convection” boundary condition leads to a black hole to total mass ratio of 0.6~0.6.
@@
Quasi-stars must remain stable long enough for the central black hole to accrete a sufficient amount of material to create a heavy seed. There are then a number of stability concerns when it comes to this sustainability, including the effects of surface winds. Dotan_2011 argued that radiation-driven winds are prominent enough to unravel the envelope on a shorter timescale than that of black hole accretion, stifling the black hole growth for all quasi-star masses below a few $\times~10^{5}$ M. ball2012quasistars employed the Reimers wind scheme, a simple analytic wind based on the mass, radius, and luminosity of the star (Reimers1975) and show that the winds have very little effect on the lifetime of the star and thus the final black hole mass. However, this scheme is likely not appropriate for the extreme conditions in supermassive or quasi-stars and was implemented as a first approximation. Fiacconi2015 suggested that taking into account a previously neglected advection term in the radiation-driven wind calculation used by Dotan_2011 significantly decreases the impact of the wind on the evaporation timescale of the quasi-star.
V MESA Implementation of Quasi-Star evolution
@@
Modules for Experiments in Stellar Astrophysics (MESA) (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023) is a collection of open-source suites designed to tackle a wide range of problems in computational stellar astrophysics. In this work, we are particularly interested in the 1-D stellar evolution module, MESA star. This module utilizes adaptive mesh refinement and advanced timestep controls to solve fully coupled structure and composition equations. MESA also uses a series of independently constructed modules to provide equation of state, opacity, nuclear reaction rates, element diffusion date, and atmosphere boundary conditions.
V.1 Quasi-stars in MESA
@@
In previous works, we developed modifications to MESA in order to model the quasi-star scenario (Campbell_2025). The black hole was implemented through adjustment of the boundary conditions and by including a sub-grid model for the black hole accretion and luminosity. Due to the massive black hole at the center of a quasi-star, general relativistic effects become important. We account for this using the Tolman-Oppenheimer-Volkov correction – a modification to the equation for hydrostatic equilibrium via a change to the gravitational constant $G$, which is recalculated within MESA at every timestep. Additionally, we use the Ledoux criterion for convection along with a mixing length of $\alpha_{mlt}=2$ as has also been used in previous models for supermassive stars (Paxton et al., 2011; Herrington_2023).
@@
To create an initial model, we evolve a typical stellar model with an initial mass of 2020 M and an envelope accretion rate of $0.1~\textrm{M}_{\odot}\textrm{yr}^{-1}$ until it reaches $10,000$ M, which is the total object mass for our fiducial quasi-star models. This method is similar to those used in Herrington_2023 when modeling supermassive stars. From this initial model, we then implement the aforementioned methods to account for the central black hole. This approach to initialization is not always necessary but does not significantly alter the quasi-star structure or lifetime in an impactful way and can thus be used regardless of necessity.
@@ V.2 Boundary Conditions
where $\rho_{0}$ is the density of the innermost cell, and $n=1$ or $3$ according to the structure of the surrounding star and whether the angular momentum is transported inward or outward, respectively (Quataert_2000). We then solve the integral
@@
V.2.1 Bondi Radius
The first inner boundary condition we adopt is the same as originally used by ball2012quasistars – the Bondi radius $R_{B}$. This is the radius at which the thermal energy of the particles equals that of the gravitational potential energy such that:
@@ V.2.1 Bondi Radius
where $m$ is the particle mass, $c_{s}$ is the sound speed, $M_{\mathrm{BH}}$ is the black hole mass and $G$ is the gravitational constant. Anything within this radius will eventually fall into the black hole, although local fluid motions within may still move both toward and away from the black hole (ball2012quasistars; Ball2012MNRAS).
V.2.2 Saturated-Convection Radius
Recently coughlin2024 proposed a new, tighter inner boundary condition for quasi-stars where convection transports energy outward at the local maximum efficiency. The result is a deeper hydrostatic region than obtained with the original Bondi conditions because the material drifts slowly inwards, compared to the free fall that the Bondi condition assumes.
@@ V.2.2 Saturated-Convection Radius
These equations use dimensionless parameters $\xi=r/r_{\rm i}$ and $m_{\rm i}=m(r)/M_{\rm i}$, where $M_{\rm i}$ is the mass contained within $r_{\rm i}$. In this case, $K_{i}$ is the only free parameter to solve saturated convection solutions, and therefore can be used to determine the inner radius given the luminosity and mass of the object (coughlin2024). Additionally, our density profile must now be $\rho\propto r^{-1/2}$ if convection is to transport all of the energy around the black hole, similar to a convection dominated accretion flow (CDAF) solution Quataert_2000. We then adjust $M_{\mathrm{cav}}$ accordingly.
V.3 Accretion & Luminosity
We now discuss the accretion schemes we have implemented within MESA for modeling quasi-stars. The defining characteristic of quasi-stars, and what makes them a viable heavy seed formation channel, is the fact that they are able to accrete at the Eddington rate of the entire object (begelman2006). This allows the seed to grow within the appropriate time constraints as set by the observation of the earliest SMBHs (Castellano_2022; Naidu_2022; Harikane_2023; Labb__2023; Natarajan2024).
We first begin with the convection-limited Bondi accretion rate, discussed previously in ball2012quasistars. This is given by:
@@ V.3 Accretion & Luminosity
Using this accretion rate and luminosity along with the Bondi radius, we obtain qualitatively similar results to those found in Part 3 of ball2012quasistars (Campbell_2025).
@@
Our fiducial models, shown in Figs. 1 and 2, begin with a central black hole mass of 10 M and a total mass of $10^{4}$ M. We used a primordial composition of 75% hydrogen and 25% helium and chose $\alpha$ values of 0.5, 0.8, 1.0, and 1.4 to explore both potential regimes where the accretion rate is suppressed (e.g., by angular momentum) and where it is enhanced (e.g., by magnetic fields).
V.4 Results
@@
We observe similar final black hole masses to those seen in Ball2012MNRAS and coughlin2024 for the Bondi radius and saturated-convection radius respectively, with a factor of $\sim 5$ increase in the saturated-convection radius over the Bondi radius for $\alpha=1$ and $1.4$. Our models then also show that these mass limits are indeed not physical, but rather a limitation of the integration scheme – the entire calculation, and thus the black hole growth, halts when the inner boundary reaches the surface of the quasi-star.
@@
Figure 2: Kippenhahn diagrams showing the profile of the quasi-star without winds in both radial (left) and mass (right) coordinates for both the Bondi (top) and saturated-convection (bottom) models. We show the Schwarzschild region (black), boundary region (grey), envelope (white), and the photosphere (black line). Both models shown use $\alpha=1$.
@@
Of note is the step-like increase in the Bondi radius. This is likely due to the connection of convective regions that were previously separated by interstitial radiative regions (ball2012quasistars). This merging would cause the star to become fully convective, sharply decreasing the sound speed, thus increasing the Bondi radius.
@@
VI Winds & Mass Loss
Up to this point, our models have only included mass loss due to the conversion of a portion of the accreted mass to photons. However, mass loss due to surface winds is far more prominent and may occur at a much higher rate (Dotan_2011). Winds in “typical” stars are fairly well understood and have a wide variety of models for various star types (Reimers1975; deJager1988; Nieuwenhuijzen1990; Bloecker1995; Nugis2000; Vink2001; vanLoon2005; Bj_rklund_2021). In this section, we explore a few of these models that are built into MESA and that best satisfy the conditions within quasi-stars as first-principle approximations. Additionally, we implement a super-Eddington radiation-driven wind scheme from Dotan_2011 that is specific to supermassive and quasi-stars.
VI.1 Reimers Scheme
Our first wind scheme is that of Reimers1975, which was also implemented in ball2012quasistars and will be referred to hereafter as the Reimers scheme. This scheme is commonly used for red giants and includes a scaling factor $\eta$ to allow for different mass loss efficiencies. The mass loss can be analytically calculated and is given by:
@@ VI.1 Reimers Scheme
where $M$, $L$, and $R$ are the total mass, luminosity, and radius in solar units, respectively, and we use a typical value $\eta=0.5$ (Reimers1975; Paxton et al., 2011).
VI.2 “Dutch” Scheme
Next we choose to explore a wind scheme used frequently for massive stars, defined as the “Dutch” scheme within MESA. This scheme is a combination of the wind models from deJager1988, vanLoon2005, and Nugis2000 where the transitions between them occur based on the surface temperature. Specifically, the combination used is based on Glebbeek2009 where they formulated the scheme to work for stars with masses $a~few\times 100$ M. For surface temperature $T_{\mathrm{eff}}>10^{4}$ K, the scheme of vanLoon2005 is used while that of deJager1988 is used for $T_{\mathrm{eff}}\leq 10^{4}$ K. It should be noted that the third model in the Dutch scheme of Nugis2000 is not used for our purposes, since it only occurs when the hydrogen mass fraction is <0.4<0.4. The temperatures and luminosities involved in the massive stars that the Dutch scheme was formulated for make this a better fit for quasi-stars, though it is likely still not appropriate. However, this scheme is built into MESA and serves as an improvement on the Reimers scheme.
@@
VI.3 Super-Eddington Radiation-Driven Wind Schemes
Our final wind schemes come from Dotan_2011 and Fiacconi2015 which were derived specifically for quasi-stars. These schemes account for the effects of the super-Eddington luminosities and the accompanying outflows. Dotan_2011 calculates approximate hydrostatic solutions for the accretion powered envelope allowing for mass loss at the surface, resulting in a mass loss rate given by:
@@ VI.3 Super-Eddington Radiation-Driven Wind Schemes
where $R_{\star}$, $\rho_{\star}$, and $c_{\star}$ are the radius, density, and sound speed at the surface, respectively. The addition of the Mach number $\mu$ allows the case where the sonic point (i.e., where the gas velocity is $c_{s}$) is not the critical point where the net forces balance each other (i.e., $R_{\star}$ when in equilibrium) (Dotan_2011; Fiacconi2015).
This scheme, however, was not modeled solving the full equations of motion and neglected the term for advection energy. Fiacconi2015 continued to numerically solve these equations with the inclusion of both diffusion and advection of energy. They found that the mass loss rate due to these winds can be reasonably well fit with power laws in the $M_{\textrm{BH}}$-M plane using:
@@
We implement this as our radiation-driven wind scheme within MESA, with a few caveats. Firstly, this model does not take into account acceleration of the material when leaving the envelope. This acceleration occurs at a radius on the length-scale of the density scale height. However, on average this scale-height is of order $1\%~R_{\star}$, so can safely be ignored for this work. This scheme also does not account for the effects of rotation, which is outside the scope of this work (Fiacconi2015).
VI.4 Results